# Suppressing Overestimation in Q-Learning through Adversarial Behaviors

### v1

@@ 5.2 Empirical Results
Grid World For exploration, we employ the $\epsilon$-greedy policy with a count-based $\epsilon(s)=1/\sqrt{n(s)}$, where $n(s)$ represents the number of visits to state $s$. The step-size is designated as $\alpha(s,a)=1/n(s,a)^{0.8}$, where $n(s,a)$ denotes the number of visits [^3] If double estimators $Q_{1}$ and $Q_{2}$ are employed with asynchronous updates, then each estimator $Q_{i}$ must be equipped with its own counter $n_{i}(s,a)$ that records the number of visits to the pair $(s,a)$ when $Q_{i}$ is updated. to the state-action pair $(s,a)$.
@@
The left subfigure in Figure 2 shows a comparative analysis of several algorithms. Using Hasselt’s reward function from Hasselt (2010), an algorithm is supposed to underestimate the value of the non-terminal transitions to reach the goal. Q-learning tends to prioritize exploration over reaching the goal due to its overestimation, which results in the average reward per step remaining around $-1$. Minmax Q-learning demonstrates superior performance compared to Q-learning, as its average reward per step tends to converge to $0.2$. With appropriate reward shifts, the minmax version of DAQ markedly enhances the performance of minmax Q-learning, as illustrated by the progression from the blue curve (minmax Q-learning) to the purple (DAQ minmax version). Here, we set $b_{1}=-5$ and $b_{2}=-10$ for both DAQs, which were determined based on some experiments in Figure 5(a) in the supplemental document.
@@
The right subfigure in Figure 2 depicts results corresponding to Wang’s reward function in Wang et al. (2020), where overestimation is more beneficial for reaching the goal state. Due to its inherent bias towards underestimation corresponding to the goal state, Double Q-learning persistently explores, showcasing the slowest convergence to the optimal average reward per step. Within the non-reward-shifting models, minmax Q-learning provides the best performance. This suggests that in certain environment setups, the minmax Q-learning holds inherent advantages over the maxmin Q-learning. The forementioned non-DAQ models all converge to suboptimal values for the averaged reward per step in their early stages of evolution. This implies that, in the scenarios with a constant step-size, they possess larger biases in their error bounds. Conversely, in the decreasing step-size scenario, this means a slower rate of convergence. However, by implementing the negative reward shifts $b_{1}=-5$ and $b_{2}=-10$, both minmax and maxmin versions of DAQ exhibit accelerated convergence to the optimal average reward per step.
Figure 2: Experiments with grid world environment, where $\gamma=0.95$ were employed. Learning curves are averaged over 10,000 experiments, and the moving averages with a window size of 100 are shown in the vivid lines. In Figure 5 in the supplement, experiments were conducted with different parameter values, $\alpha=0.1$ or $\epsilon=0.1$, where the trends were similar.
@@
Sutton’s Example [^4] The experimental results with $\mu=+0.1$ is shown in Appendix A.2 in the supplement, and the result is not much different. with $\mu=-0.1$ Since $\mu$ is less than 00, the optimal action at state $\mathsf{A}$ is $\mathsf{right}$. Consequently, 5% is the optimal ratio, illustrated as the black horizontal lines. Figure 3 depicts the learning curves of several algorithms. The first, second, and last subfigures display results associated with $8$, $20$, and $100$ actions at state $\mathsf{B}$, respectively. On the first subfigure with $8$ actions, with a constant step-size $\alpha=0.1$, all the discussed algorithms converge to the optimal ratio. However, DAQs attain precisely a 5% ratio relatively early on, around the $180$-th episode, outperforming Double Q-learning which continues to exhibit biases at this stage.
@@
The final subfigure depicts the learning curves associated with $100$ actions. As expected, minmax Q-learning learns at a slower pace; therefore, we have omitted its graph for clarity. When negative shifts are implemented in the reward function, the related DAQ versions converge to the optimal 5% ratio more swiftly compared to both Q-learning and maxmin Q-learning but at a more gradual pace than the previous cases. To expedite convergence, synchronous DAQ was also tested. A notable enhancement in the learning curve is evident transitioning from the purple curve to the olive one. In the synchronized DAQ, it learns to take $\mathsf{right}$ action quickly until around $250$-th episode. We additionally conducted tests with decreasing $\alpha_{n}=\frac{10}{n+100}$, and the learning curves closely resemble those in the constant step-size case, detailed in Figure 9 in the supplement.
Figure 3: Experiments with Sutton’s example with $\mu=-0.1$, experimented with $\epsilon=0.1$, $\alpha=0.1$ and $\gamma=1$. The three subfigures show the learning curves for different numbers of actions at state $\mathsf{B}$, where the learning curves were averaged over 10,000 experiments. For DAQs, $b_{1}=-1$ and $b_{2}=-2$ were used. In each subfigures, the number of episodes are different in order to show the convergence of the algorithms.
Weng’s Example In this experiment, we utilize $M\in\{8,20,100\}$, with results displayed in Figure 4. Note that in Sutton’s Example, the action space enlarges, whereas in Weng’s Example, it’s the state space that expands as $M$ increases. Since every action from state $\mathsf{1}$ to $\mathsf{M}$ results in a negative expected reward, the optimal action at state $\mathsf{0}$ is the $\mathsf{right}$ action. Given $\epsilon$-greedy exploration, a 5% ratio of taking the $\mathsf{left}$ action is optimal, represented by 5% horizontal lines. The learning curves of DAQs are highlighted with pink and purple colors, revealing that DAQ versions attain the optimal action faster compared to the other algorithms. In DAQ, $b_{1}=+0.0001$ and $b_{2}=+0.0002$ are incorporated into the reward function to attain better performance; these are minor compared to the reward $\mathcal{N}(-0.1,1)$ or 00. When the shifts get closer to zero, the DAQs converge faster as shown in Figure 9(a) in the supplemental document. Note that the graphs for DAQ versions begin at nearly 50%, in contrast to others starting at around 40%. With the incorporation of shifted rewards, DAQ versions achieve more uniform exploration, preventing biased exploration.
Figure 4: Experiments with Weng’s example, where $\alpha_{n}=\frac{10}{n+100}$, $\epsilon=0.1$ and $\gamma=1$ are used. The three subfigures correspond to the results with different numbers of states $M$. The learning curves were averaged over 10,000 experiments. Similar results can be obtained with the constant step-size $\alpha=0.1$ as shown in Figure 9(b) in the supplement.
6 Conclusion
@@ Appendix A Additional Empirical Results
Here, we give more experimental results which are omitted from Section 5.2.
A.1 Grid World
@@
The experimental results using the Grid World environment is depicted in this subsection.
@@
In Figure 5, parameter values other than mainly suggested in Section 5.2 are tested. The first column is the same as Figure 2 for comparison, and the second and the last column use parameter values of $\alpha=0.1$ or $\epsilon=0.1$. Figure 4(a) shows the result using Hasselt’s reward function. The second subfigure presents a comparative analysis featuring a decreasing step-size and $\epsilon$-greedy exploration with a constant $\epsilon=0.1$, and the result is not much different from the first subfigure. The last subfigure provides a comparative analysis employing a constant step-size $\alpha=0.1$ and a constant $\epsilon=0.1$, where the DAQs continue to display superior results. Figure 4(b) shows the result using Wang’s reward function. Both minmax Q-learning and maxmin Q-learning show significantly better results than Q-learning and Double Q-learning. In the second subfigure, the distinctions between DAQs and other considered algorithms become clearer when a constant step-size $\alpha=0.1$ is used. Employing a constant $\epsilon=0.1$ does not significantly alter the comparative results, as is shown in the last subfigure.
In Figure 6, shift values other than mainly suggested in Section 5.2 are tested. Figure 5(a) shows the result using Hasselt’s rewards. For both DAQs, we empirically checked that the negative value of $b_{i}$ facilitates converging to optimal while the positive values hinder convergence. Performance is better with $(b_{1},b_{2})=(-5,-10)$ than $(b_{1},b_{2})=(-1,-2)$, and even slightly better with $(b_{1},b_{2})=(-20,-30)$. However, we decided to use $-5$ and $-10$ in Figure 2 since the difference is not much distinguishable. Figure 5(b) shows the results using Wang’s rewards, especially when $(b_{1},b_{2})=(-1,-2)$ is used for DAQs at Grid World Experiment using Wang’s reward function. The result with decreasing step-size is not much different from that of Figure 2, but the convergence at constant step-size was better at Figure 2 which used $(b_{1},b_{2})=(-5,-10)$.
@@
(a) Result using Hasselt’s reward function. Decreasing step-size is used on the first and the second subfigures, and count-based $\epsilon(s)$ is also used on the first subfigure.
@@
(b) Result using Wang’s reward function. Decreasing step-size is used on the first subfigure, and count-based $\epsilon(s)$ is also used on the first and the second subfigures.
Figure 5: Result of Grid World Experiment. $\gamma=0.95$ is used. Data is averaged over 10,000 experiments, and the moving average with a window size of 100 is shown in vivid lines. The horizontal lines at $0.2$ represent the average reward per step for optimal policy.
@@
(a) Result using Hasselt’s reward function with various shift values, which are presented in titles of each subfigures. $\alpha=1/n(s,a)^{0.8}$ and count-based $\epsilon(s)$ are used. The graphs of Q-learning, Double Q-learning, minmax Q-learning and maxmin Q-learning are the same for all three subfigures for better comparison, and only their moving averages are drawn here.
@@
(b) Result with Wang’s reward function. $\alpha=1/n(s,a)^{0.8}$ is used on the first subfigure, and count-based $\epsilon(s)$ is also used on the first and the second.
Figure 6: Result of Grid World Experiments. $\gamma=0.95$ is used. Data is averaged over 10,000 experiments, and the moving average with a window size of 100 is shown in vivid lines. The horizontal lines at $0.2$ represent the average reward per step for optimal policy.
A.2 Sutton’s Example
Sutton’s Example with $\mu=+0.1$ In this scenario, actions at state $\mathsf{B}$ are anticipated to yield positive rewards, thus making $\mathsf{left}$ the optimal action at state $\mathsf{A}$. To explore this environment, the $\epsilon$-greedy exploration with $\epsilon=0.1$ is employed. Consequently, after the convergence of the Q iterate to $Q^{*}$, it is optimal to take the action $\mathsf{left}$ at state $\mathsf{A}$ 95% of the time, aligning with the logic of $\epsilon$-greedy exploration. The action $\mathsf{right}$ is chosen for the remaining 5%. These two ratios are illustrated by the black horizontal lines in Figure 7. Each curve other than the lines represents the evolution of the ratio of selecting $\mathsf{left}$ actions. In this setting, the designer has the autonomy to determine the number of available actions at state $\mathsf{B}$, which, in this instance, has been set to eight.
Figure 6(a) illustrates the learning curves associated with the decreasing step-size, denoted as $\alpha_{n}=\frac{10}{n+100}$ for episode $n$. Minmax Q-learning reaches the optimal ratio considerably quicker than the other algorithms, and both Q-learning and maxmin Q-learning approach the optimum but exhibit a preference for the $\mathsf{right}$ action during the exploration phase. The grey curve in the graph represents the maxmin version of DAQ with $b_{1}=+0.01$ and $b_{2}=+0.02$, which were determined from experiments in Figure 8. A significant alteration in the learning curve can be observed when comparing the grey and the orange curves. Even though $0.01$ or $0.02$ are relatively small quantities when compared to the rewards, these reward shifts significantly aid the algorithm in learning an optimal policy effectively and efficiently.
Figure 6(b) displays the learning curves associated with a constant step-size, $\alpha=0.1$. One can observe that only the DAQ versions converge to the optimal ratio, while the remaining algorithms settle at values with larger biases. DAQs with positive $b_{i}$ maintain exploratory actions up to approximately the $300$-th episode, after which they acquire the proficiency to choose the optimal $\mathsf{left}$ action.
(a) Experiments using decreasing step-sizes, where minmax Q-learning exhibits the best performance. Minmax Q-learning can be considered as minmax DAQ with $b_{i}=0$.
(b) Experiments with constant step-sizes, where $b_{1}=+1$ and $b_{2}=+2$ are used to both DAQs. Note that the curves of the two DAQs are almost overlapped.
Figure 7: Experiments with Sutton’s example with $\mu=+0.1$, where 95% is optimal, $\epsilon=0.1$ and $\gamma=1$ were used. The two subfigures show learning curves with different step-size rules. The learning curves were averaged over 5,000 experiments.
Figure 8 is another experiment with various shift values tested to Sutton’s Example with $\mu=+0.1$, and the grounds for the chosen shift values at Figure 6(a). Starting from positive $(b_{1},b_{2})=(+1,+2)$ and negative $(b_{1},b_{2})=(-1,-2)$, we adjusted the amount by 10 times, from $10^{-3}$ to $10^{1}$ scale. All positive shifts disturb the algorithm to converge: the larger the value, the slower the convergence. On the other hand, among the negative values, $10^{-2}$ is the most effective while $10^{-3}$ converges faster to the optimal at first, but is soon saturated and shows the same slope with Q-learning or maxmin Q-learning. Especially, positive shifts are shown in Figure 8. $(b_{1},b_{2})=(+1,+2)$ or $(+10,+20)$ cannot learn any specific policy, and the algorithm starts to converge as the amount of shift decreases.
Figure 8: Result of various shift values, which are presented in titles of each subfigures, using Sutton’s Example with $\mu=+0.1$. Here, 95% line is optimal and 5% is sub-optimal. $\alpha_{n}=\frac{10}{n+100}$, $\epsilon=0.1$ and $\gamma=1$ are used. Data is averaged over 5,000 experiments. The graphs of Q-learning, Double Q-learning, minmax Q-learning and maxmin Q-learning are the same for all three subfigures for better comparison.
@@ A.2 Sutton’s Example
Figure 9 shows the results using Sutton’s Example with $\mu=-0.1$. The graphs are similar to Figure 3, except for the complete convergence of Q-learning and maxmin Q-learning here. Decreasing step-size makes the algorithms converge to exact optimal compared to constant step-size. We additionally suggested the graphs of synchronized DAQ for all three subfigures. When the size of the action space is small, the synchronized DAQ fluctuates a lot, though steadily gets closer to the optimal and achieves the optimal faster than any other algorithms.
@@
Figure 9: Result using Sutton’s Example with $\mu=-0.1$, where a 5% line is optimal. From the first to the last subfigure: 8 actions, 20 actions, and 100 actions from state $\mathsf{B}$, respectively. $\epsilon=0.1$, $\alpha_{n}=\frac{10}{n+100}$ and $\gamma=1$ are used. Data is averaged over 10,000 experiments. $b_{1}=-1$ and $b_{2}=-2$ are used for shifting. We increased the number of episodes from 600 to 1,000 and 4,000 to check the convergence of the algorithms.
A.3 Weng’s Example

### v3

@@ V-B Empirical Results
Grid World We adopted the exploration settings from the work of [3]: $\epsilon$-greedy behavior policy with a count-based $\epsilon(s)=1/n(s)^{0.5}$ and step-size $\alpha(s,a)=1/n(s,a)^{0.8}$, where $n(s)$ represents the number of visits to state $s$ and $n(s,a)$ denotes the number of visits [^3] If double estimators $Q_{1}$ and $Q_{2}$ are employed with asynchronous updates, then each estimator $Q_{i}$ must be equipped with its own counter $n_{i}(s,a)$ that records the number of visits to the pair $(s,a)$ when $Q_{i}$ is updated. to the state-action pair $(s,a)$. Discount factor $\gamma$ is set to $0.95$. Results with different $\alpha$ or $\epsilon$ settings are shown in Fig. 9 in the supplement.
@@
The left subfigure in Fig. 4 shows a comparative analysis of several algorithms. Using Hasselt’s reward function $r^{H}_{t}$, an algorithm is supposed to underestimate the value of the non-terminal transitions to reach the goal. Q-learning tends to prioritize exploration over reaching the goal due to its overestimation, which results in the average reward per step remaining around $-1$. Minmax Q-learning demonstrates superior performance compared to Q-learning, as its average reward per step tends to converge to $0.2$. With appropriate reward shifts, the minmax DAQ markedly enhances the performance of minmax Q-learning. Here, we set $b_{1}=-5$ and $b_{2}=-10$ for both DAQs, which were determined based on some experiments in Fig. 9 in the supplemental document.
@@
The right subfigure in Fig. 4 depicts results corresponding to Wang’s reward function $r^{W}_{t}$, where overestimation is more beneficial for reaching the goal state. Due to its inherent bias towards underestimation corresponding to the goal state, double Q-learning persistently explores, showcasing the slowest convergence to the optimal average reward per step. Within the non-reward-shifting models, minmax Q-learning provides the best performance. This suggests that in certain environment setups, the minmax Q-learning holds inherent advantages over the maxmin Q-learning. The forementioned non-DAQ models all converge to suboptimal values for the averaged reward per step in their early stages of evolution. This implies that, in the scenarios with a constant step-size, they possess larger biases in their error bounds. Conversely, in the decreasing step-size scenario, this means a slower rate of convergence. However, by implementing the negative reward shifts $b_{1}=-5$ and $b_{2}=-10$, both minmax and maxmin DAQ exhibit accelerated convergence to the optimal average reward per step.
Figure 4: Average rewards from grid world environment. DAQs achieve optimal policy for both reward functions. The moving averages with a window size of 100 are shown in the vivid lines.
Figure 5: Experiments with Sutton’s MDP with $\mu=-0.1$. DAQs highly outperform other algorithms. In each subfigure, the number of episodes are different in order to show the convergence of the algorithms. For DAQs, $b_{1}=-1$ and $b_{2}=-2$ were used.
Sutton’s MDP [^4] We here show the case of $\mu=-0.1$, and the other case, $\mu=+0.1$ whose result is not much different, is shown in Appendix A-B in the supplement. with $\mu=-0.1$ We adopted the parameter settings in [16]: $\epsilon=0.1$, $\alpha=0.1$ and $\gamma=1$. We additionally conducted tests with decreasing step-size and the learning curves closely resemble those in the constant step-size case, detailed in Fig. 11 in the supplement.
@@
Since $\mu$ is less than 00, the optimal action at state $\mathsf{A}$ is $\mathsf{right}$. Consequently, 5% is the optimal ratio, illustrated as the black horizontal lines. Fig. 5 depicts the learning curves of several algorithms. The first, second, and last subfigures display results associated with $8$, $20$, and $100$ actions at state $\mathsf{B}$, respectively. On the first subfigure with $8$ actions, all the discussed algorithms converge to the optimal ratio. However, DAQs attain precisely a 5% ratio relatively early on, around the $180$-th episode, outperforming double Q-learning which continues to exhibit biases at this stage.
@@
The final subfigure depicts the learning curves associated with $100$ actions. As expected, minmax Q-learning learns at a slower pace while the synchronized DAQ learns optimal policy quickly. Detailed analysis is shown in Appendix B.
Weng’s MDP We adopted the exploration setting in [21]: $\epsilon$-greedy behavior policy with $\epsilon=0.1$ and step-size $\alpha_{n}=\frac{10}{n+100}$ for episode $n$. Discount factor $\gamma$ is set to $1$. In this experiment, we utilize $M\in\{8,20,100\}$, with results displayed in Fig. 6. Note that in Sutton’s MDP, the action space enlarges, whereas in Weng’s MDP, it’s the state space that expands as $M$ increases. Since every action from state $\mathsf{1}$ to $\mathsf{M}$ results in a negative expected reward, the optimal action at state $\mathsf{0}$ is the $\mathsf{right}$ action. Given $\epsilon$-greedy exploration, a 5% ratio of taking the $\mathsf{left}$ action is optimal, represented by 5% horizontal lines. The learning curves of DAQs are highlighted with pink and purple colors, revealing that DAQ versions attain the optimal action faster compared to the other algorithms. In DAQ, $b_{1}=+0.0001$ and $b_{2}=+0.0002$ are incorporated into the reward function to attain better performance; these are minor compared to the reward $\mathcal{N}(-0.1,1)$ or 00. When the shifts get closer to zero, the DAQs converge faster as shown in Fig. 15 in the supplemental document. Note that the graphs for DAQ versions begin at nearly 50%, in contrast to others starting at around 40%. With the incorporation of shifted rewards, DAQ versions achieve more uniform exploration, preventing biased exploration.
Figure 6: Experiments with Weng’s MDP. The three subfigures correspond to the results with different numbers of states $M$. Similar results can be obtained with the constant step-size $\alpha=0.1$ as shown in Fig. 16 in the supplement.
VI Conclusion
@@ Appendix A Additional Empirical Results
Here, we give more experimental results which are omitted from Section V-B.
A-A Grid World
@@
The experimental results using the grid world environment are depicted in this subsection. Results are averaged over 10K experiments with $\gamma=0.95$. In Section V-B, results with $\epsilon(s)=1/n(s)^{0.5}$, $\alpha(s,a)=1/n(s,a)^{0.8}$, $b_{1}=-5$ and $b_{2}=-10$ have shown in Fig. 4.
@@
In Fig. 7 and Fig. 8, parameter values other than mainly suggested in Section V-B are tested. Average rewards from grid world environment are shown, and the moving average with a window size of 100 is shown in vivid lines. The horizontal lines at $0.2$ represent the average reward per step for optimal policy. The graphs from Fig. 4 are repeated in the first column for comparison, and the second and the last column use parameter values of $\alpha=0.1$ or $\epsilon=0.1$. Fig. 7 shows the result using $r^{H}_{t}$. The second subfigure presents a comparative analysis featuring a decreasing step-size and $\epsilon$-greedy exploration with a constant $\epsilon=0.1$, and the result is not much different from the first subfigure. The last subfigure provides a comparative analysis employing a constant step-size $\alpha=0.1$ and a constant $\epsilon=0.1$, where the DAQs continue to display superior results. Fig. 8 shows the result using $r^{W}_{t}$. Both minmax Q-learning and maxmin Q-learning show significantly better results than Q-learning and double Q-learning. In the second subfigure, the distinctions between DAQs and other considered algorithms become clearer when a constant step-size $\alpha=0.1$ is used. Employing a constant $\epsilon=0.1$ does not significantly alter the comparative results, as is shown in the last subfigure.
@@
Figure 7: Result using $r^{H}_{t}$. Decreasing step-size is used on the first and the second subfigures, and count-based $\epsilon(s)$ is also used on the first subfigure.
@@
Figure 8: Result using $r^{W}_{t}$. Decreasing step-size is used on the first subfigure, and count-based $\epsilon(s)$ is also used on the first and the second subfigures.
In Fig. 9 and Fig. 10, shift values other than mainly suggested in Section V-B are tested. Average rewards from grid world environment are shown, and the moving average with a window size of 100 is shown in vivid lines. The horizontal lines at $0.2$ represent the average reward per step for optimal policy. Fig. 9 shows the result using $r^{H}_{t}$. For both DAQs, we empirically checked that the negative value of $b_{i}$ facilitates converging to optimal while the positive values hinder convergence. Performance is better with $(b_{1},b_{2})=(-5,-10)$ than $(b_{1},b_{2})=(-1,-2)$, and even slightly better with $(b_{1},b_{2})=(-20,-30)$. However, we decided to use $-5$ and $-10$ in Fig. 4 since the difference is not much distinguishable. Fig. 10 shows the results using $r^{W}_{t}$, especially when $(b_{1},b_{2})=(-1,-2)$ is used for DAQs. The result with decreasing step-size is not much different from that of Fig. 4, but the convergence at constant step-size was better at Fig. 4 which used $(b_{1},b_{2})=(-5,-10)$.
@@
Figure 9: Result using $r^{H}_{t}$ with various shift values, which are presented in titles of each subfigures. $\alpha(s,a)=1/n(s,a)^{0.8}$ and count-based $\epsilon(s)$ are used. The graphs of minmax Q-learning and maxmin Q-learning are the same for all three subfigures for better comparison, and only their moving averages are drawn here.
@@
Figure 10: Result using $r^{W}_{t}$ with shift $b_{1}=-1$ and $b_{2}=-2$. $\alpha=1/n(s,a)^{0.8}$ is used on the first subfigure, and count-based $\epsilon(s)$ is also used on the first and the second.
A-B Sutton’s MDP
The experimental results using Sutton’s MDP with $\epsilon=0.1$ and $\gamma=1$ are depicted in this subsection. In Section V-B, results with $\mu=-0.1$ and $\alpha=0.1$ have shown in Fig. 5. The third subfigure of Fig. 5 depicts the learning curves associated with $100$ actions. As expected, minmax Q-learning learns at a slower pace; therefore, we have omitted its graph for clarity. When negative shifts are implemented in the reward function, the related DAQ versions converge to the optimal 5% ratio more swiftly compared to both Q-learning and maxmin Q-learning but at a more gradual pace than the previous cases. To expedite convergence, synchronous DAQ was also tested. A notable enhancement in the learning curve is evident transitioning from the purple curve to the olive one. In the synchronized DAQ, it learns to take $\mathsf{right}$ action quickly until around $250$-th episode.
@@ A-B Sutton’s MDP
Sutton’s MDP with $\mu=-0.1$ Fig. 11 shows the results using decreasing step-size denoted as $\alpha_{n}=\frac{10}{n+100}$ for episode $n$. The graphs are similar to Fig. 5, except for the complete convergence of Q-learning and maxmin Q-learning here. Decreasing step-size makes the algorithms converge to exact optimal compared to constant step-size. We additionally suggested the graphs of synchronized DAQ for all three subfigures. When the size of the action space is small, the synchronized DAQ fluctuates a lot, though steadily gets closer to the optimal and achieves the optimal faster than any other algorithms.
@@
Figure 11: Result using Sutton’s MDP with $\mu=-0.1$ and decreasing step-size, where a 5% line is optimal. From the first to the last subfigure: 8 actions, 20 actions, and 100 actions from state $\mathsf{B}$, respectively. $b_{1}=-1$ and $b_{2}=-2$ are used for shifting. We increased the number of episodes from 600 to 1,000 and 4,000 to check the convergence of the algorithms. Results are averaged over 10K runs.
Sutton’s MDP with $\mu=+0.1$ In this scenario, actions at state $\mathsf{B}$ are anticipated to yield positive rewards, thus making $\mathsf{left}$ the optimal action at state $\mathsf{A}$. To explore this environment, the $\epsilon$-greedy exploration with $\epsilon=0.1$ is employed. Consequently, after the convergence of the $Q$ iterate to $Q^{*}$, it is optimal to take the action $\mathsf{left}$ at state $\mathsf{A}$ 95% of the time, aligning with the logic of $\epsilon$-greedy exploration. The action $\mathsf{right}$ is chosen for the remaining 5%. These two ratios are illustrated by the black horizontal lines in Fig. 12 and Fig. 13. These two figures show learning curves with different step-size rules, with Sutton’s MDP with $\mu=+0.1$, where 95% is optimal. The learning curves were averaged over 5K runs. Each curve other than the lines represents the evolution of the ratio of selecting $\mathsf{left}$ actions. In this setting, the designer has the autonomy to determine the number of available actions at state $\mathsf{B}$, which, in this instance, has been set to eight.
Fig. 12 illustrates the learning curves associated with the decreasing step-size $\alpha_{n}=\frac{10}{n+100}$. Minmax Q-learning reaches the optimal ratio considerably quicker than the other algorithms, and both Q-learning and maxmin Q-learning approach the optimum but exhibit a preference for the $\mathsf{right}$ action during the exploration phase. The grey curve in the graph represents the maxmin version of DAQ with $b_{1}=+0.01$ and $b_{2}=+0.02$, which were determined from experiments in Fig. 14. A significant alteration in the learning curve can be observed when comparing the grey and the orange curves. Even though $0.01$ or $0.02$ are relatively small quantities when compared to the rewards, these reward shifts significantly aid the algorithm in learning an optimal policy effectively and efficiently.
Fig. 13 displays the learning curves associated with a constant step-size, $\alpha=0.1$. One can observe that only the DAQ versions converge to the optimal ratio, while the remaining algorithms settle at values with larger biases. DAQs with positive $b_{i}$ maintain exploratory actions up to approximately the $300$-th episode, after which they acquire the proficiency to choose the optimal $\mathsf{left}$ action.
Fig. 14 is another experiment with various shift values tested to Sutton’s MDP with $\mu=+0.1$, and the grounds for the chosen shift values at Fig. 12. Starting from positive $(b_{1},b_{2})=(+1,+2)$ and negative $(b_{1},b_{2})=(-1,-2)$, we adjusted the amount by 10 times, from $10^{-3}$ to $10^{1}$ scale. All positive shifts disturb the algorithm to converge: the larger the value, the slower the convergence. On the other hand, among the negative values, $10^{-2}$ is the most effective while $10^{-3}$ converges faster to the optimal at first, but is soon saturated and shows the same slope with Q-learning or maxmin Q-learning. Especially, positive shifts are shown in Fig. 14. $(b_{1},b_{2})=(+1,+2)$ or $(+10,+20)$ cannot learn any specific policy, and the algorithm starts to converge as the amount of shift decreases.
Figure 12: Experiments using decreasing step-sizes, where minmax Q-learning exhibits the best performance. Minmax Q-learning can be considered as minmax DAQ with $b_{i}=0$.
Figure 13: Experiments with constant step-sizes, where $b_{1}=+1$ and $b_{2}=+2$ are used to both DAQs. Note that the curves of the two DAQs are almost overlapped.
Figure 14: Result of various shift values using Sutton’s MDP with $\mu=+0.1$ and $\alpha_{n}=\frac{10}{n+100}$. Here, 95% line is optimal and 5% is sub-optimal. Data is averaged over 5K runs. The graphs of Q-learning, double Q-learning, minmax Q-learning and maxmin Q-learning are the same for all three subfigures for better comparison.
A-C Weng’s MDP
The experimental results using Weng’s MDP are depicted in this subsection. Results are averaged over 10K experiments with $\epsilon=0.1$ and $\gamma=1$. In Section V-B, results with $\alpha_{n}=\frac{10}{n+100}$ have shown in Fig. 6.